OpenStudy (anonymous):

Find the measure of A to the nearest whole degree. Use the table for sine. a is similar to

OpenStudy (anonymous):



OpenStudy (anonymous):

sin-1 (4/20)= 11.53 so about 12 degrees

OpenStudy (anonymous):

just plug it in the calculator

OpenStudy (anonymous):

thankyou. would you mind helping me with about 9 more questions

OpenStudy (anonymous):

sure if i can :D

OpenStudy (anonymous):

are you going to ask the questions?

OpenStudy (anonymous):

ok..Find the measure of A to the nearest whole degree. Use the table for sine

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

same thing inverse sin which is sin-1(7/10)=44.43

OpenStudy (anonymous):

so it's 44 degrees

OpenStudy (anonymous):

Find the measure of A to the nearest whole degree. Use the table for sine

OpenStudy (anonymous):

do you have a calculator?

OpenStudy (anonymous):

yea a ti 85

OpenStudy (anonymous):

well can't you do it yourself now? it doesn't take a second

OpenStudy (anonymous):

it's 19 anyway

OpenStudy (anonymous):

i need help like doing what i'm doing. im only in the 9th grade and ive never learned about this

OpenStudy (anonymous):

you should have a button in yous calculator that says sin-1

OpenStudy (anonymous):

you push it and then divide the opposite side of the angle by the hypotenuse

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

ok and i should get 19

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

try it and see if it works

OpenStudy (anonymous):

it didn't work

OpenStudy (anonymous):

did you find the sin button?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

above it there's a sin-1 right?

OpenStudy (anonymous):

not on mine i have the TI 84.... but when press the 2nd button and sin it says Sin-1 right

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

so click 2nd sin then put ( opposite side of A divided by the hypotenuse)

OpenStudy (anonymous):

what would that be in this equation

OpenStudy (anonymous):

so sin-1(3/9)= 19

OpenStudy (anonymous):

the opposite side is the side not touching the angle and the hypotenuse is the longest side

OpenStudy (anonymous):

3/9 try it again now

OpenStudy (anonymous):

i keep getting .3271946968

OpenStudy (anonymous):

did you click sin-1(3/9) exactly like that sin-1(3/9)

OpenStudy (anonymous):

if it doesn't work do sin-1 .3333333333 wich is 3/9

OpenStudy (anonymous):

don't push anything between sin-1 and the (3/9)